# Beyond-$\Lambda$CDM Cosmologies with Bayesian Neural Networks II: Massive Neutrinos, Baryonic Feedback, and the Theoretical Error

### v1

@@ Abstract
We study the capacity of Bayesian Neural Networks (BNNs) to detect new physics in the dark matter power spectrum. As in previous studies, the Bayesian Cosmological Network (BaCoN) classifies spectra into one of 5 classes: $\Lambda$CDM, $f(R)$, $w$CDM, Dvali-Gabadaze-Porrati (DGP) gravity and a “random” class, with this work extending it to include the effects of massive neutrinos and baryonic feedback. We further develop the treatment of theoretical errors in BaCoN-II, investigating several approaches and identifying the one that best allows the trained network to generalise to other power spectrum modelling prescriptions. In particular, we compare power spectra data produced by EuclidEmulator2, HMcode and halofit, all supplemented with the halo model reaction to model beyond-$\Lambda$CDM physics. We investigate BNN-classifiers trained on these sets of spectra, adding in Stage-IV survey noise and various theoretical error models. Using our optimal theoretical error model, our fiducial classifier achieves a total classification accuracy of $\sim$ 95% when it is trained on EuclidEmulator2-based spectra with modification parameters drawn from a Gaussian distribution centred around $\Lambda$CDM ($f(R)$$\sigma_{fR0}=10^{-5.5}$, DGP: $\sigma_{r\mathrm{c}}=0.173$, $w$CDM: $\sigma_{w0}=0.097$, $\sigma_{wa}=0.32$). This strengthens the promise of this method to glean the maximal amount of unbiased gravitational and cosmological information from forthcoming Stage-IV galaxy surveys.
@@
pubyear: 2024pagerange: Beyond-$\Lambda$CDM Cosmologies with Bayesian Neural Networks II: Massive Neutrinos, Baryonic Feedback, and the Theoretical ErrorB
@@ 1.2 Machine Learning
A deep neural network (DNN) provides a highly nonlinear map between the input (in this case a cosmological quantity) and the output (in this case a classification of theoretical model). DNNs have shown to be successful in many classification problems, predominantly outside physics (Mehta et al., 2019; Lin et al., 2017; LeCun et al., 2015). Bayesian Neural Networks (BNNs) try to model the uncertainty by replacing each weight in the network by a distribution initialised to a prior (MacKay, 1992; Neal, 1996; Blundell et al., 2015; Gal & Ghahramani, 2016; Charnock et al., 2022; Valentin Jospin et al., 2020). In a previous paper (Mancarella et al., 2020), the authors suggested the employment of BNNs to supplement cosmological MCMC analyses. This works by classifying an input power spectrum using the BNN, and then using that classification to select various degrees of freedom or apply stronger priors on model parameters, both of which can be applied to heighten efficiency of subsequent MCMC analyses.
@@ 1.3 Overview
This paper is laid out as follows. Section 2 presents our theoretical and numerical setup. Section 3 then describes the data preparation before it can be passed to the BNN, along with a discussion of the BNN’s architecture. Section 4 presents our results and discusses the performance of the BNN. Section 5 presents our conclusions and outlines future plans.
@@ 2.1 Bayesian and Convolutional Neural Networks
The architecture and machine learning techniques used here are the same as in our previous work. We point the reader to Mancarella et al. (2020) for the details.
@@ 3.2 Noise Model
3.3 Neural Network
We train the BNN with the noisy training data. The shape of the data can be adapted at this stage, e.g. with scale cuts. We employ a training data set of 20,000 spectra per class (of which 15% are used as a validation set) to train our most accurate models. The architecture of the BNN, hyperparameters, loss function etc. remain unchanged from the initial BaCoN and are described in detail in Mancarella et al. (2020). We train the network for 50 epochs or until the training accuracy has become stationary. This is the most time consuming part of the machine learning process. The list of data sets used for training and testing is displayed in Table 2 and has been made publicly available here.
Once we have trained the network, we test it with the noisy test data. As in Mancarella et al. (2020), every example requires a classification probability of 50% or above to be considered as classified. Otherwise we count it as not classified (N.C.). After passing all 1000 test data spectra through the network with 10 noise realisations each, we obtain the confusion matrix for this test set. We will present the test results for many different models and tests in the next section.
@@ 3.3 Neural Network
We will present the theoretical error model, results for our baseline model and the influence of the blue marked components in Figure 1 in the next section.

### v2

# and the Theoretical Error

@@ Abstract
We study the capacity of Bayesian Neural Networks (BNNs) to detect new physics in the dark matter power spectrum. As in previous studies, the Bayesian Cosmological Network (BaCoN) classifies spectra into one of 5 classes: $\Lambda$CDM, $f(R)$, $w$CDM, Dvali-Gabadaze-Porrati (DGP) gravity and a ‘random’ class, with this work extending it to include the effects of massive neutrinos and baryonic feedback. We further develop the treatment of theoretical errors in BaCoN-II, investigating several approaches and identifying the one that best allows the trained network to generalise to other power spectrum modelling prescriptions. In particular, we compare power spectra data produced by EuclidEmulator2, HMcode and halofit, all supplemented with the halo model reaction to model beyond-$\Lambda$CDM physics. We investigate BNN-classifiers trained on these sets of spectra, adding in Stage-IV survey noise and various theoretical error models. Using our optimal theoretical error model, our fiducial classifier achieves a total classification accuracy of $\sim$ 95% when it is trained on EuclidEmulator2-based spectra with modification parameters drawn from a Gaussian distribution centred around $\Lambda$CDM ($f(R)$$\sigma_{fR0}=10^{-5.5}$, DGP: $\sigma_{r\mathrm{c}}=0.173$, $w$CDM: $\sigma_{w0}=0.097$, $\sigma_{wa}=0.32$). This strengthens the promise of this method to glean the maximal amount of unbiased gravitational and cosmological information from forthcoming Stage-IV galaxy surveys.
@@
pubyear: 2024pagerange: Classifying Modified Gravity and Dark Energy Theories with Bayesian Neural Networks: Massive Neutrinos, Baryonic Feedback, and the Theoretical ErrorB
@@ 1.2 Machine Learning
A deep neural network (DNN) provides a highly nonlinear map between the input (in this case a cosmological quantity) and the output (in this case a classification of theoretical model). DNNs have shown to be successful in many classification problems, predominantly outside physics (Mehta et al., 2019; Lin et al., 2017; LeCun et al., 2015). Bayesian Neural Networks (BNNs) try to model the uncertainty by replacing each weight in the network by a distribution initialised to a prior (MacKay, 1992; Neal, 1996; Blundell et al., 2015; Gal & Ghahramani, 2016; Charnock et al., 2022; Valentin Jospin et al., 2020). In a previous paper (Mancarella et al., 2022), the authors suggested the employment of BNNs to supplement cosmological MCMC analyses. This works by classifying an input power spectrum using the BNN, and then using that classification to select various degrees of freedom or apply stronger priors on model parameters, both of which can be applied to heighten efficiency of subsequent MCMC analyses.
@@
One major advantage of using a Bayesian instead of a Deterministic Neural Network is the treatment of classification uncertainties. The uncertainty in the classification by a neural network can be divided into the intrinsic uncertainty of the data and the uncertainty due to the machine learning model itself. These components are known as aleatoric and epistemic uncertainties, respectively (Kiureghian & Ditlevsen, 2009). Our approach to deal with uncertainties is threefold. First, we introduce a fifth ‘random’ class that accounts for out of distribution samples. It represents cosmologies that are not represented by the cosmological models chosen for the other training data sets. It flags previously unseen cosmological classes and might also get activated when the cosmological parameters of the test samples are far outside of the prior range of the training data. We investigate the effectiveness of this ‘random’ class in Section 5. Second, we marginalise over the weight distribution in the Bayesian layers. This is achieved by averaging over multiple passes through the network, each drawing different Monte Carlo (MC) samples from the weight distributions. At the same time, we can use the distribution of output probabilities to construct an estimate for the uncertainty of the classification result. The computation of the distribution of MC averages is presented in Section 3.5.
However, this does not erase all of the epistemic uncertainty of the network. The training data is not a perfect representation of real observations. We try to compensate this with the third part of our uncertainty treatment, by introducing a theoretical error in the training process. A detailed discussion can be found in Section 4.1. While this can account for the biggest part of the systematic error of the training data, some of the code-specific features of the training data will remain and bias the neural network. We investigate the error induced by different power spectrum prescriptions in Section 4.4. In the next section we will have a look at other implementations of neural networks for the analysis of modified gravity and dark energy models.
1.3 Neural Networks for Beyond-$\Lambda$CDM models
When testing the application of machine learning techniques to modified gravity scenarios, the well studied $f(R)$ class is particularly popular. It’s Hu-Sawicki parameterisation (Hu & Sawicki, 2007) has been implemented in many simulation suites.
García-Farieta et al. (2024) use real-space density fields and power spectra of $f(R)$ DM-only simulations as training data for their network. It is constructed from different Bayesian modifications of the ResNet18 architecture – a convolutional neural network with 18 layers. The main goal of their work is to constrain cosmological parameters from $f(R)$ simulations. The Bayesian networks perform well with exception for the estimation of the MG parameter itself. They adopt multiplicative normalising flows for the approximate posteriors of the BNN parameters. This method of BNN training was tested in their previous work (Hortúa et al., 2023) for the estimation of $\Lambda$CDM parameters from N-body simulations. An interesting combination of a BNN and a Recurrent Neural Network are used in Escamilla-Rivera et al. (2020) to investigate dark energy models. This study is not based on LSS data but uses the distance modulus from supernovae data. They implement a Bayesian network to obtain uncertainty contours.
Moving away from BNNs, Ocampo et al. (2024b) use a deterministic neural network to classify $f(R)$ spectra based on the evolution of $f\sigma_{8}$ values. Their focus is on the interpretability of the chosen network architecture, e.g. studying feature importance. They follow a similar approach in Ocampo et al. (2024a) but this time using Cosmic Microwave Background data. Another test of a $f(R)$-vs-$\Lambda$CDM classification has been performed in Peel et al. (2019). They use a deterministic convolutional neural network to classify simulated weak-lensing convergence maps and specifically test the influence of massive neutrinos. They find that adding noise significantly lowers the network’s performance.
Apart from neural networks, many other machine learning approaches have been used to study alternatives to $\Lambda$CDM. Gangopadhyay et al. (2023) use a genetic algorithm to construct a more model-independent test of dark energy theories. A similar approach has been taken in Arjona & Nesseris (2020) to search for deviations from $\Lambda$CDM in the background expansion. A genetic algorithm can be employed to optimise the design of neural network architectures, as it was shown in Gómez-Vargas et al. (2023) for the equation of state of a quintessence model.
It seems that to our best knowledge not many multi-class MG classifiers have been constructed. Most likely this is the case because the risk of degeneracies increases when more MG models are considered at the same time. BaCoN-II tries to circumvent this problem by using LSS data from highly nonlinear scales. This requires accurate models for physics affecting nonlinear scales like massive neutrinos and baryonic feedback.
@@ 1.4 Overview
This paper is laid out as follows. Section 2 presents our theoretical and numerical setup, along with a discussion of the BNN’s architecture. Section 3 then describes the data preparation, the training and testing process as well as an analysis of the calibration and uncertainty of the trained network. Section 4 presents our results and discusses the performance of the BNN. Section 5 presents our conclusions and outlines future plans.
@@ 2.1 Bayesian and Convolutional Neural Networks
The architecture and machine learning techniques used here are described in Section 2.4, they are adopted from our previous work in Mancarella et al. (2022).
@@ 2.3 Data Sets
2.4 Network Architecture and Training Methods
The key question behind BaCoN is whether we can design a neural network that can recognise features of different cosmological models in the matter power spectrum and distinguish them from $\Lambda$CDM. This requires 3 main ingredients: layers that perform the feature analysis of the data, components that regularise the training and ensure a well-calibrated model, and a mathematical model for the weight distributions and their optimisation in the training process.
Figure 1: Architecture of BaCoN-II: Schematic representation of network layers for the classification of a power spectrum with 100 $k$-bins and 4 $z$-bins into 5 classes. The output dimension for each layer is noted on the left side. Additional details for the 1D-Convolutional layers (filter number, kernel size and stride) and for the 1D-MaxPooling layers (pooling size and stride) are shown on the right side. Schematic created with Netron.
[…]
Apart from our image-analysing layer we need methods that reduce the dimensionality and renormalise the layer’s output. We use 1D max pooling layers and batch normalisation layers for this, respectively. A global average pooling layer is used to flatten the feature maps into a dense layer with 32 nodes. Finally, a softmax layer with 5 nodes produces the classification output.
In total, the architecture of BaCoN-II, shown in Figure 1, consists of 6605 trainable parameters. They are tuned by passing training data through the network and minimising a loss function with an optimisation algorithm. A backpropagation algorithm finds the minimum of the derivative of the loss to update the parameters and match the network’s output closer to the training labels. This method is called gradient descent. Here, we are using the adam optimisation algorithm (Kingma & Ba, 2014). We also have to set a learning rate (lr) that dictates the size of the steps in the direction of the gradient. It changes after every full pass of all training samples through the network, called an epoch $e$. For $N_{\mathrm{T}}$ samples in the training set we get the learning rate
$\text{lr}(e)=0.01\times 0.95^{(e/N_{\mathrm{T}})}\,.$ (5)
In a BNN, the single values for weights in a traditional deterministic layer are replaced by parameterised distributions (MacKay, 1992; Neal, 1996). The goal of training the network is to find the posterior distributions $p(w|\mathcal{D})$ of the network’s weights $w$ given the training data $\mathcal{D}$ starting from a prior distribution of the weights $p(w)$. A common way to find an approximation for the posterior weight distributions is variational inference (Jordan et al., 1999; Graves, 2011; Charnock et al., 2020). It uses a family of distributions $q_{\theta}(w)$ parameterised by the parameter $\theta$ that are closely matched to the posterior $p(w|\mathcal{D})$ in the training process. One way to achieve this is by minimising the Kullback-Leibler divergence term (KL) (Kullback & Leibler, 1951). Our loss function is constructed from a negative log-likelihood term in combination with a KL-divergence between the variational distribution and the prior that acts as a regularisation term.
Unfortunately, computing the gradients for variational inference is difficult. A reparameterisation trick (Kingma & Welling, 2013; Kingma et al., 2015) can be used to perform backpropagation but then the gradients of a batch become correlated. Wen et al. (2018) have developed the Flipout technique to decorrelate the weights in a mini-batch. Flipout has become widely popular for the use in BNNs and has been implemented in TensorFlow Probability (Dillon et al., 2017). We are using the implementation of the Flipout technique in TensorFlow-Probability-0.14.0 in all convolutional and dense layers pictured in Figure 1. In comparison with other established BNN training methods, Flipout has been shown to be one of the higher performing variational inference methods (Hortúa et al., 2020). Other works have implemented alternatives to the Gaussian variational posterior $q_{\theta}(w)$ for a further improved calibration (Hortúa et al., 2023). Therefore, it is important to test whether the network architecture in combination with the training data leads to a well-calibrated network. We will present the results of the calibration test for BaCoN-II in Section 3.4.
To evaluate the training progress we are using a dedicated validation set. After every epoch we compute the classification accuracy for the spectra in the validation set to see how the training progresses. The training is finished when the validation accuracy converges to a stable configuration and the loss stabilises at a minimum. The next section will explain the details of the practical procedures for training and testing the network.
@@ 3.2 Noise Model
3.3 Training and Testing the Neural Network
We train the BNN with the noisy training data. The shape of the data can be adapted at this stage, e.g. with scale cuts. We employ a training data set of 20,000 spectra per class (of which 15% are used as a validation set) to train our most accurate models. The architecture of the BNN, hyperparameters, loss function etc. remain unchanged from the initial BaCoN by Mancarella et al. (2022) and are described in detail in Section 2.4. We train the network for 50 epochs or until the training accuracy has become stationary. This is the most time consuming part of the machine learning process. The list of data sets used for training and testing is displayed in Table 2 and has been made publicly available here.
Once we have trained the network, we test it with the noisy test data. As in Mancarella et al. (2022), every example requires a classification probability of 50% or above to be considered as classified. Otherwise we count it as not classified (N.C.). In total we use 1000 test data spectra and pass them through the network with 10 noise realisations each. To make full use of the stochastic features of BaCoN-II, we take the average of 100 MC samples per every noisy sample for the confusion matrix. So we technically pass every noisy spectrum 100 times through the network and sample from the weight distributions. The final classification result for every test spectrum is an average of the classification outputs from all MC samples. Hence, the confusion matrix is a generalised representation of the performance of the trained network marginalised over the weights. When rerunning the test code for the same trained model and test data, the average over the accuracy of 100 MC samples changes overall by less than 1%. Therefore, we do not state errors in the confusion matrices.
@@ 3.3 Training and Testing the Neural Network
We will present the theoretical error model, results for our baseline model and the influence of the blue marked components in Figure 2 in Section 4.
3.4 Calibration
The final output of the softmax layer of the network sums to one and can be interpreted as probabilities. For this to be an reasonable interpretation, we need to ensure that the network is well calibrated. Especially since research on DNNs has suggested that modern networks tend to have miscalibration issues (Guo et al., 2017). We construct a calibration diagram of our baseline model to measure the reliability of the network’s output. First, we run a test of our baseline model with the same test data and testing procedure as in our main result shown in Figure 8. For consistency, we add 10 noise realisations to every test spectrum and the $\mu$ values are generated from 100 MC samples. Then, we bin the predictions $\mu_{i}$ for every class $i$ and compute the average $\hat{\mu}(b,i)$ in every bin $b$. We compare this to the true probability $P_{\mathrm{true}}(b,i)$ which is the frequency of the positive label for this class $i$ for the spectra in bin $b$. The calibration of every individual class is displayed in Figure 4 following the design by Hollemans (2020). A perfectly calibrated network would produce a diagonal line in the calibration diagram. We can see that most classes look well calibrated with the exception of $f(R)$ and the random class that show multiple bins that are strongly underconfident. This is actually not a calibration effect but a result of the low number of spectra in these bins caused by the high confidence for this classes and the lack of spectra with $\mu$ far from 0 or 1. Therefore, we’ve added a histogram with the bin counts for every class and mark bins with a fainter colour if they contain less than 0.5 % of the total number of test spectra.
Figure 4: Calibration of the baseline model: The bins represent the difference between the predicted probability $\mu_{i}$ of a class $i$ compared to the true probability given by the True-False number ratio of test spectra for this class in this $\mu$ bin.The Expected Calibration Error (ECE) defined in Equation 7 is displayed for every class individually. The lower row shows the relative bin count (fraction of test spectra per bin to the total test size). If the bin contains less than 0.5 % of the total number of test spectra it is marked with a fainter colour in the calibration diagram as it barely influences the overall calibration. The classifications are made with the baseline model and baseline test data set (see Figure 8 for the full test result).
The population of bins is automatically accounted for in the Static Calibration Error (SCE), a measure for the overall miscalibration of a multiclass network. The SCE of BNNs is computed in the same way as for deterministic networks, with the only difference that we average $\mu$ over 100 MC samples for every spectra in the Bayesian network. We can calculate it as an averaged sum over the Expected Calibration Errors (ECE) of the individual classes following Nixon et al. (2019) as
SCE $=\frac{1}{N}\sum_{i=1}^{N}\sum_{b=1}^{B_{i}}\frac{n_{bi}}{N_{\text{tot}}}\left|P_{\mathrm{true}}(b,i)-\hat{\mu}(b,i)\right|$ (7)
$=\frac{1}{N}\sum_{i=1}^{N}\mathrm{ECE}_{i}\,,$
with the total number of classes $N$, the number of bins per class $B_{i}$ and the total number of test samples $N_{\mathrm{tot}}$. The ECEs for every class are shown in Figure 4. The individual ECEs are all below 2 % and combined give in a total SCE for BaCoN-II of 1.1 %. This shows that the network is overall well calibrated. Hence, we can reasonably use the output $\mu$ as an estimate for the probability. We will have a more detailed look at the properties and uncertainties of this output in the next section.
3.5 Uncertainty Estimates
We now present a test for the uncertainty of the BaCoN-II classification. This provides an insight into the internal workings of the Bayesian network and reflects the advantages of using a Bayesian instead of a deterministic neural network. First, we need to understand the statistics of the classification in a Bayesian network. The output of the final softmax layer is a vector with one component $p_{i}$ for each class $i$ of a total of $N$ classes. [^1] In this work we train BaCoN-II for 5 classes, so $N=5$, but this can be adapted easily. As they add to 1, the values of $p_{i}$ can be interpreted as the probabilities for each class, when the network is calibrated correctly (see Figure 4). In a deterministic neural network we only get a single prediction for $p_{i}$, as the weights are fixed and repeating the analysis will result in the same output vector. Hence, we have no way to tell whether the classification result is reliable. A BNN on the other hand provides an intrinsic method to compute the uncertainty of the classification. When passing a spectrum through the trained network we Monte-Carlo sample from the distribution of the weights $w$. Multiple passes will result in different output probabilities $p_{i,\alpha}$, where we use the greek index $\alpha$ to indicate the MC sample. We can marginalise over the weights of the network by averaging over a number of $N_{\mathrm{MC}}$ MC samples. This gives an approximate value for the average output probability of the softmax layer
$\mu_{i}\approx\frac{1}{N_{\mathrm{MC}}}\sum_{\alpha=1}^{N_{\mathrm{MC}}}\,p_{i,\alpha}\,.$ (8)
We classify the example spectrum by assigning the label of the class with the highest MC average $\mu_{i}$, if it exceeds a set threshold probability $p_{\mathrm{th}}$. To get a confidence estimate, we need the probability distribution of these MC averages $\mu$, which we denote as $\mathcal{F}$. We construct this function by modelling a distribution for a random variable $x_{i}$ (representing the softmax output of the network) with mean $\mu_{i}$ and the full covariance of the classification $\Sigma$. The covariance of a trained network is computed following Kwon et al. (2020) and Mancarella et al. (2022) as
$\Sigma=\frac{1}{N_{\mathrm{MC}}}\sum_{\alpha=1}^{N_{\mathrm{MC}}}\,\left(\text{diag}(p_{\alpha})-p_{\alpha}^{\otimes 2}\right)+\frac{1}{N_{\mathrm{MC}}}\sum_{\alpha=1}^{N_{\mathrm{MC}}}(p_{\alpha}-\mu)^{\otimes 2}\,,$ (9)
with $y^{\otimes 2}=yy^{T}$. The first term of the covariance represents the aleatoric uncertainty and the second term models the epistemic uncertainty. The resulting output probability distribution is then given by
$\mathcal{F}(x;\mu,\Sigma)$ $=\,\delta\Big{(}1-\sum_{j=1}^{N}x_{j}\Big{)}\times\sqrt{N}$ (10)
×$\times\prod_{i=1}^{N-1}{\mathcal{\tilde{N}}}\Big{(}\left[{B}^{-1}(x-\mu)\right]_{i};0,\left[{B}^{-1}\Sigma{B}\right]_{ii}\Big{)}\,,$
where $\mathcal{\tilde{N}}$ is a multivariate Gaussian truncated between 0 and 1 and $B$ is the matrix that diagonalises $\Sigma$. The derivation of this form of $\mathcal{F}$ and how to construct it from the distribution of MC-sample probabilities is described in the appendix of Mancarella et al. (2022). Finally, to obtain a classification statement and label of the spectrum we need to impose a threshold probability $p_{\mathrm{th}}$, in our case $50\,\%$. The probability $P_{i}$ of labelling the spectrum as a class $i$ is determined from the fraction of samples that fulfil the criterion $x_{i}>p_{\mathrm{th}}$. This gives an estimate of the uncertainty for the classification result.
Figure 5 displays the probability distribution $\mathcal{F}$ for a specific spectrum with one fixed noise realisation. We are adding the same noise that we are adding to test data, i.e. cosmic variance at large scales. The spectrum is passed through the network 100 times. We use these 100 MC samples to compute $\mu$ and $\Sigma$ and finally construct $\mathcal{F}$. The plots are created by drawing 1000 samples from the final distribution. We show the label probabilities $P_{i}$ for each class as well as $P_{\mathrm{unclassified}}$, when the output does not exceed $p_{\mathrm{th}}=0.5$ for any of the classes.
Figure 5 shows $\mathcal{F}$ for four different spectra – two are $f(R)$ and two are $w$CDM examples. We choose these two classes because, as we will later see in Figure 8, $w$CDM shows a higher degeneracy with other classes while $f(R)$ gets classified very confidentially. The two spectra have class parameters that produce a strong deviation from $\Lambda$CDM. They are classified correctly $100\,\%$ of the time with an output probability of $\mu_{i}=1$ for the correct class throughout. This changes drastically when we look at spectra that have only a small modification compared to $\Lambda$CDM. The $\mu_{i}$ distributions widen and we see higher chances of unclassified and wrongly classified results. This trend shows how a difficult detection due to small features leads to a low confidence of the classification that can be captured by plotting $\mathcal{F}$ and the label probabilities $P_{i}$. It is a measure for the uncertainty of the classification while marginalising over the weight distributions of the network. The $w$CDM spectrum in particular shows a high degeneracy with DGP, which we will discuss in more detail in Section 4.2. The distribution of $\mathcal{F}$ as plotted in Figure 5 is a powerful tool to estimate the uncertainty for the classification of one specific spectrum. This is relevant for observational data where we only have a single spectrum with one noise realisation. An estimate of confidence in the classification of observational data is an essential component for the method to be useful. Further, if the spectrum is determined to be ’Not Classified’, $\mathcal{F}$ will indicate if there are any classes that can be excluded safely and which classes to focus further analysis on.
Figure 5: Output probability distribution $\mathcal{F}$ for some example spectra: $x_{i}$ represents the output of the network, that can be interpreted as the probability for the respective class $i$. $\mathcal{F}$ is the probability distribution of the MC averages for one specific power spectrum with a fixed noise realisation (cosmic variance only). $\mathcal{F}$ is computed with Equation 10 based on the average $\mu_{i}$ and variance $\Sigma$ expected for 100 MC samples. We draw 1000 samples from $\mathcal{F}$ to display the distribution here. The classifications are made with the baseline EE2-based model (see Figure 8) for two $f(R)$ and two $w$CDM spectra with less and more extreme class-specific parameters.